can u list me all  formulae with descriptions  for the topic mechanical of fluids?

Grade:11

1 Answers

Ajay Kumar gupta
7 years ago
1. Pressure=atmospheric pressure +height×density×g2. Density=mass÷volume3. Upthrust=decrease in weight of solid inside liquid=weight of fluid displaced by solid4. Equation of continuity Area of cross-section×velocity=remain constant(Of single tube) - AV=av5. Density+1/2×density×(velocity)*2+density×g×height=constant6. Velocity of efflux=2×g×height below the water surface upto hole"Time taken by liquid to reach ground={(2×total height of liquid-2×height below the water surface upto hole)*1/2}÷(g)*1/2"Horizontal displacement within this time is RS range=velocity of efflux × time taken by liquid to reach the ground